OpenStudy (anonymous):

To ship books, a store charges a fixed fee for the carton and an additional amount for each book. It costs $30 to ship 10 books and $35 to ship 12 books. What is the fixed fee in dollars?

OpenStudy (anonymous):

Alrighty lets make this problem something a little more workable... Lets make it a linear equation.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Lets have the coordinates (10,30) and (12,35) as the points. (35-30)/(12-10) = 5/2 = 2.5 as a slope.

OpenStudy (anonymous):

Fixed rate would be 2.5.